OpenStudy (anonymous):

29=b/-4+15 ( b over -4 )

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

\[29=\frac{b}{-4}+15\rightarrow 29-15=\frac{b}{-4}\rightarrow -4(29-15)=b \rightarrow b=-56\]

OpenStudy (anonymous):

b=-56

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

whats the wow for?

OpenStudy (anonymous):

thats a lot of numbers hahaha:)

OpenStudy (anonymous):

lol...it's all good. I just broke it down instead of giving you the answer, this way you can understand the process

OpenStudy (anonymous):

haha that works. Can you help me with some more?

OpenStudy (anonymous):

sure but you have 5 minutes... I have meeting at 7:15

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

2.7=1.3-2d

OpenStudy (anonymous):

\[2.7=1.3-2d \rightarrow2.7-1.3=-2d \rightarrow \frac{2.7-1.3}{-2}=d \rightarrow d=-.7\]

OpenStudy (anonymous):

wow i was waaay off. Dx

OpenStudy (anonymous):

thaaanks:D